# Characterizing the MrDeepFakes Sexual Deepfake Marketplace

### v2

@@ 8 Conclusion
The MrDeepFakes Platform. MrDeepFakes’ terms of service prohibit using site content for commercial endeavors, redistributing content, and sending users unsolicited emailsmessages, but they do not prohibit scraping or automated data collectionotherwise prohibit scraping. The site’s robots.txt only disallows crawling by the Internet Archive (User-agent: ia_archiver). All of the information we scraped from the website scraped information is accessible to anyone with an account.   [^4] Account creation is free and does not require e-mail verification; site operators have also stated that users are not obligated to provide “real” information. (Account creation is free and does not require e-mail verification; site operators have also stated that users are not obligated to provide “real” information.) We acknowledge that in scraping the website, we may have increased server load and bandwidth on the part of the platform; however, we minimized the number of requests made to the platform, and our request volume is relatively low compared to the organic activity of tens of thousands of accounts that the platform services. We stress that we , such that our study has minimal impact on the platform.
We did not collect or store any video or image data, only collecting relevant metadata required to answer our research questions.
@@
We did consider We considered whether to publish the name of the platform itself in this paper, as similar prior work opted to not reveal the name of the platforms in question platform names as to not draw potential attention to its their existence [tseng20tools]. However, we decided to name the MrDeepFakes platform as it has been named directly in prior work [timmerman2023studying, popova2020reading] and is the top search result for celebrity sexual deepfakes. We hope that naming this platform can enable future research and potential regulation on the issue of further research and policy on sexual deepfake creation and mainstreaming.
Users of the platform.
In examining user data, we collected only publicly available data, and
@@
Users of MrDeepFakes . In examining public user profiles, the only potentially personally identifying information we collected was the account username and the site-assigned user ID. Accounts are pseudonymous: there are no fields for “real” names. We never attempted to deanonymize any userin our dataset , and we did not interact with any community members in any fashion community members (e.g., via direct messages or public posts). In our report, we take careful note to not include any users’care to exclude any user’s account details when reporting quotes and opt to only report report only aggregate statistics to protect user privacy.
Targets of sexual deepfakes.
To avoid drawing undue attention to any targets mentioned in the forum, we redact and paraphase any personally identifying information in the quotes we reported and elect not to name any target in the textPer a publication requirement from the USENIX Security program committee, we paraphrased quotes to mitigate the possible re-identification of quoted users.
Researchers.
@@
Researchers. Given that this paper required qualitative analysis of text-based forum data, the researchers involved in these tasks were directly exposed to potentially disturbing content that might cause significant disturbing text discussion and deepfake content that could cause psychological or emotional harms. distress. The research team was aware of these potential harms but recognized that exposure to this content was unavoidable to gain domain knowledge, understand the affordances of rendered webpages, and verify collected data. We ultimately made an informed decision to move forward with our study because we believed in the timeliness and importance of investigating this ecosystem. To mitigate these issuesharms, the researchers had regular check-ins regarding their psychological well being, well-being, worked together as a team to process MrDeepFakes content, had access to therapy and counseling services, and took breaks as needed.
@@
We are committed to sharing our data and code with other researchers at academic, non-commercial institutions seeking to conduct future work. However, given the sensitive nature of the data and disturbing nature of the contentdata, we will not make this data available publicly. not make this data available publicly. We note that the MrDeepFakes site content is available on the open Internet and that motivated actors can easily access the contentfor themselves. However, we do not want to further enable malicious actors seeking to use MrDeepFakes data to potentially harm others. We are committed to sharing our data and our codebooks with the Artifact Evaluation committee to ensure our artifacts meet the USENIX Open Science standards.
Appendix A MrDeepFakes Forum mapSection Map
@@ Appendix B File Sharing Platforms
Appendix C Investigated Deepfake production guides we examinedProduction Guides
@@
Appendix D Breakdown of sexual Sexual deepfake video-related video discussiontopics
Total: The breakdown, by number of threads and posts, of forum discussion topics relating to sexual deepfake videos. The total number of threads is 6,259threads , comprising 33,496 posts in the Adult DeepFake Videos (NSFW / 18+) and DeepFake Video Creation Tools sections.
@@ Appendix E Paid Request Codebook
Table 5: Codebook for need-finding interviews Codebook for Paid Deepfake Requests —Codebook for categorizing the different themes that emerged from our need-finding interviews with eight journaliststhematic analysis of a sample (N=263) of paid requests for deepfake videos.

### v3

@@ Acknowledgments
The MrDeepFakes Platform. MrDeepFakes’ terms of service prohibit using site content for commercial endeavors, redistributing content, and sending users unsolicited messages, but they do not otherwise prohibit scraping. The site’s robots.txt only disallows crawling by the Internet Archive (User-agent: ia_archiver). All scraped information is accessible to anyone with an account. (Account creation is free and does not require e-mail verification; site operators have also stated that users are not obligated to provide “real” information.) We acknowledge that in scraping the website, we increased server load and bandwidth on the part of the platform; however, we minimized the number of requests made to the platform, and our request volume is relatively low compared to the organic activity of tens of thousands of accounts that the platform services, such that our study has minimal impact on the platform.
@@
We did not collect any video or image data, only relevant metadata required to answer our research questions. We considered whether to publish the name of the platform in this paper, as similar prior work opted to not reveal platform names as to not draw potential attention to their existence [63]. However, we decided to name MrDeepFakes as it has been named directly in prior work [62, 52] and is the top search result for celebrity sexual deepfakes. We hope that naming this platform can enable further research and policy on sexual deepfake creation and mainstreaming.
@@
Users of MrDeepFakes. In examining public user profiles, the only potentially personally identifying information was the account username and site-assigned user ID. Accounts are pseudonymous: there are no fields for “real” names. We never attempted to deanonymize any user, and we did not interact with community members (e.g., via direct messages or public posts). In our report, we take care to exclude any user’s account details when reporting quotes and opt to report only aggregate statistics to protect user privacy. Per a publication requirement from the USENIX Security program committee, we paraphrased quotes to mitigate the possible re-identification of quoted users.
@@
Researchers. Given that this paper required qualitative analysis of text-based forum data, the researchers involved in these tasks were directly exposed to disturbing text discussion and deepfake content that could cause psychological or emotional distress. The research team was aware of these potential harms but recognized that exposure to this content was unavoidable to gain domain knowledge, understand the affordances of rendered webpages, and verify collected data. We ultimately made an informed decision to move forward with our study because we believed in the timeliness and importance of investigating this ecosystem. To mitigate these harms, the researchers had regular check-ins regarding their psychological well-being, worked together as a team to process MrDeepFakes content, had access to therapy and counseling services, and took breaks as needed.
@@
We are committed to sharing our data and code with researchers at academic, non-commercial institutions seeking to conduct future work. We note that MrDeepFakes site content is available on the open Internet and that motivated actors can easily access the content. However, we do not want facilitate malicious abuse of MrDeepFakes data to potentially harm others. Therefore, given the sensitive and disturbing nature of the data, rather than public distribution, we will selectively provide access to the resource repository upon manual verification of the requesting parties. We believe that making the data available to the research community with appropriate vetting (i.e., of affiliation, intended research purposes) balances adding friction for abuse with providing utility to the community.
Appendix A MrDeepFakes Forum Section Map
@@ Appendix B File Sharing Platforms
Appendix C Investigated Deepfake Production Guides
@@
Appendix D Sexual deepfake video discussion
The breakdown, by number of threads and posts, of forum discussion topics relating to sexual deepfake videos. The total number of threads is 6,259, comprising 33,496 posts in the Adult DeepFake Videos (NSFW / 18+) and DeepFake Video Creation Tools sections.
@@ Appendix E Paid Request Codebook
Table 5: Codebook for Paid Deepfake Requests—Codebook for categorizing the different themes that emerged from our thematic analysis of a sample (N=263) of paid requests for deepfake videos.